The length of a screw produced by a machine is normally distributed with a mean of 0.75 inches and a standard deviation of 0.01 inches. What percent of screws are between 0.72 and 0.78 inches?
Answer by stanbon(75887)

The length of a screw produced by a machine is normally distributed with a mean of 0.75 inches and a standard deviation of 0.01 inches. What percent of screws are between 0.72 and 0.78 inches?
------------------------------------
z(0.72) = (0.72-0.75)/0.01 = -0.03/0.01 = -3
----
z(0.78) = (0.78-0.75)/0.01 = 3
-----
P(-3< z <3) = normalcdf(-3,3) = 0.9973 = 99.73%
----------------------
Cheers,
Stan H.
----------------